# Orders in Chaos: Enhancing Large-Scale MoE LLM Serving with Data Movement Forecasting

### v1

@@
Abstract.
Large Language Models (LLMs) with Mixture of Experts (MoE) architectures achieve remarkable performance improvements, but their random expert selection mechanism introduces significant data movement overhead that becomes the dominant bottleneck in multi-unit serving systems.
To forecast the patterns underlying this data movement, we conduct comprehensive data-movement-centric profiling across three state-of-the-art large-scale MoE models (200B-671B) using over 24,000 requests spanning diverse workloads. With the resulting 150GB+ trace files, we perform systematic analysis from both temporal and spatial perspectives and distill six key insights to guide the design of diverse future serving systems. Taking wafer-scale GPUs as a case study, we demonstrate that minor architectural modifications leveraging our insights achieve substantial performance gains, delivering 6.3× and 4.0× average speedups on DeepSeek V3 and Qwen3, respectively. Our work provides the first comprehensive data-centric analysis of MoE models at scale. Our profiling traces and analysis results are publicly available at {https://huggingface.co/datasets/core12345/MoE_expert_selection_trace. We will also release our simulation framework shortly to facilitate future research in this area.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
Table 1. Overview of State-of-the-art MoE LLMs. They are becoming larger and sparser, incurring more data movement.
Year Month Param
Top
Experts
Select
Experts
Sparsity
Ratio
Mixtral8x7B (Mistral, 2023) 2023 12 45B 8 2 4
DBRX (Databricks, 2024) 2024 3 132B 16 4 4
Mixtral8x22B (Mistral, 2024) 2024 4 141B 8 2 4
@@ 1. Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across diverse domains, including programming assistance (Nam et al., 2024; Wang et al., 2021), translation (Achiam et al., 2023; Lu et al., 2024), and chatbots (Dam et al., 2024; Vakayil et al., 2024). To further improve model quality with limited computation cost, the Mixture of Experts (MoE) architecture has emerged as a promising path for scaling LLMs.
Unlike traditional LLMs, MoE models introduce significant data movement overhead that dominates serving latency. This overhead stems from random expert selection, which dynamically routes each request to a subset of experts instead of using all model parameters. As MoE models continue to scale with increasing parameters and expert counts (Table 1), this overhead becomes increasingly severe. Even serving small models like Mixtral 8×7B on small systems (2-4 GPUs) already sees data movement consuming over 50% of end-to-end time (Zhang et al., 2025; Go and Mahajan, 2025). Recent large-scale models like DeepSeek V3, with 32× more experts and 15× more parameters than Mixtral 8×7B, require larger systems (multiple nodes, 32+ GPUs) where data movement becomes even more frequent and costly. Thus, identifying and exploiting the underlying data movement patterns in MoE models presents a fruitful opportunity for better efficiency.
@@
If data movement in MoE models were fully unpredictable, it would present significant challenges for deployments on multi-unit systems. From a temporal perspective, the explosive growth in expert combinations would make it impossible to prefetch, cache, or replicate experts in advance. For example, large-scale MoE models like DeepSeek V3 have $C_{256}^{8}=$4,426,165,368$\$ combinations in expert selection, and such unpredictability would result in data movement like expert migrations across units, incurring substantial overhead, as inter-unit communication becomes the primary bottleneck. From a spatial perspective, if expert selection were truly random, it would lead to severe workload imbalance across units. When queries from diverse tasks are served concurrently, the number of queries assigned to each expert would vary dramatically, creating significant workload disparities. Consequently, most units would remain idle while waiting for heavily loaded units to complete processing, resulting in poor hardware resource utilization.
@@
Fortunately, as we will later show in the paper, MoE expert selections indeed have predictability that designers can exploit to reduce data movement. To uncover the inherent patterns in MoE models, we conduct a comprehensive data-movement-centric profiling of three state-of-the-art MoE models ranging from 235B to 671B parameters. As highlighted in Table 1, we profile DeepSeek V3, Llama4, and Qwen3 across 24,000 requests involving varied tasks, topics, and languages, which consumes ¿ 2000 GPU hours in total. We then collect the expert selection trace of all layers and tokens in each request to create an expert selection database of over 150 GB JSON files. From these extensive traces, we conduct a comprehensive analysis to uncover data movement patterns from both temporal and spatial perspectives, making our findings system-agnostic and applicable to various serving architectures at any scale. Through this large-scale analysis, we distill six key insights that serve as a solid foundation to our understanding of MoE data movement and directly inform future MoE LLM serving system design, addressing critical questions that have remained unanswered in the field, such as: Is there a correlation between previously selected experts and those selected later? Are there discernible rules underlying the observed expert selection skewness? Do different tasks tend to activate different experts? Our work represents the first systematic effort to characterize data movement patterns at the scale of ¿200B model across a wide range of tasks, providing actionable insights that can guide the design of next-generation MoE serving systems.
@@
  • We propose a comprehensive and systematic data-movement-centric profiling across three models between 235B and 671B to uncover the data movement patterns from both temporal and spatial perspectives.
@@
  • We collect over 70,000 expert selection traces across multiple models and datasets, totaling over 150 GB in JSON format, and will open-source all traces along with our multi-chiplet simulator to facilitate future research on MoE serving systems and architectures.
Figure 1. Inference process of MoE LLMs and the categorization method for our proposed data-centric profiling approach.
2. Background
2.1. LLM and MoE Model Architecture
Most state-of-the-art LLMs adopt a decoder-only transformer architecture that follows a token-by-token auto-regressive workflow (Vaswani et al., 2017). As shown in Figure 1(a), after users input queries, the serving process is divided into two stages: the prefill stage and the decode stage. During the prefill stage, all input tokens are processed simultaneously to generate the first output token. The decode stage follows immediately, where tokens are generated sequentially. The generated token from each iteration is appended to the input sequence to produce the next token in the following iteration.
The Mixture of Experts (MoE) mechanism is a state-of-the-art approach to improve LLM performance and has become prevalent among existing and future top-tier LLMs (Masoudnia and Ebrahimpour, 2014). As shown in Figure 1(b), MoE-based LLMs replace the feed-forward network (FFN) operations in traditional LLMs with MoE operations. In each layer, multiple experts are deployed, and each request is routed to a small subset of the most suitable experts based on a gating mechanism. This innovation enables MoE models to scale model parameters without incurring additional inference overhead, since only a fraction of parameters are activated for each request. However, this mechanism also introduces dynamic randomness, as expert selection remains unknown until gating is completed, posing new challenges for serving systems.
2.2. Prior MoE Serving Systems
Existing research on MoE serving optimization has developed numerous system-level solutions to handle data movement and target different performance and cost objectives. Edge systems like MoE-Lightning (Cao et al., 2025a) and CoServe (Suo et al., 2025) employ CPU memory offloading techniques to address GPU memory capacity constraints, while server systems such as Comet (Zhang et al., 2025) and MegaScale-Infer (Zhu et al., 2025) leverage multi-GPU architectures and address GPU-GPU communications in MoE for higher throughput. Novel hardware architectures like Duplex (Yun et al., 2024) explore processing-in-memory to accelerate data movement in MoE LLMs.
@@ 2.2. Prior MoE Serving Systems
Figure 2. Layer-level temporal correlation heatmaps for (a) Deepseek and (b) Qwen, together with (c) statistical results across all layers, reveal a strong correlation between expert selection in layer NN and that in layer $N+1$.
Figure 3. Token-level temporal correlation heatmaps for (a) Deepseek, (b) Llama, and (c) Qwen, together with (d) statistical results across all layers, demonstrate a strong correlation between the expert selection for consecutive tokens $t$ and $t+1$.
Figure 4. Expert activation patterns remain consistent across prefill and decode stages for both (a, b) cross-layer heatmap and (c, d) cross-token heatmaps. Spearman’s ratio quantified in (e, f) confirms this relation ($\geq 0.7$ means a strong relation).
3. MoE Profiling and System Insights
In this section, we conduct a data-movement-centric profiling of the expert selection behavior in three state-of-the-art MoE models: Deepseek V3, Llama4-128E, and Qwen3-235B. All results are averaged over more than 24,000 requests.
3.1. Categorization Methodology
We categorize MoE expert selection profiling results into two fundamental perspectives, as shown in Figure 1: temporal and spatial relations. These perspectives reveal distinct expert selection behaviors and motivate corresponding system strategies for mitigating data movement challenges.
@@ 3.2.1. Layer-Level
The heatmaps reveal clear cross-layer correlations with white dots highlighting specific expert pairs with significantly higher selection probabilities across adjacent layers. However, correlation patterns vary across layers within the same model and differ between models due to architectural variations. For instance, patterns between layers 3-4 differ from those between layers 30-31. Qwen3’s notably brighter heatmap indicates stronger cross-layer correlations than Deepseek. Beyond the white dots, there are also consistent bright vertical lines observed, suggesting certain experts are frequently chosen regardless of previous layer selections. These patterns indicate generally popular experts, analyzed further in subsubsection 3.3.1.
To quantify these relationships, we analyze expert pair frequencies across multiple requests and all model layers. The cumulative percentage of the most frequently selected expert pairs is shown in Figure 2(c). The results demonstrate that the top 20% of expert pairs account for 45%, 68%, and 80% of all activations in Deepseek, Qwen, and Llama4, respectively. This reveals strong, model-dependent cross-layer correlations, with Llama4 exhibiting the strongest effect and Deepseek the weakest.
Note that Llama4’s interleaved architecture places traditional FFN layers between MoE layers, so we analyze correlations between adjacent MoE layers (layer $N$ and $N+2$).
3.2.2. Token-Level
@@
Similar to layer-level patterns, cross-token heatmaps exhibit white dots, bright vertical lines, and variation across layers and models, indicating correlations between adjacent tokens. However, token-level relations reveal a distinctive pattern appearing across all models: the bright diagonal line that indicates the tendency to select the same expert across adjacent tokens. This diagonal pattern emerges predominantly in higher layers (17 and 43) but not in lower layers (1 and 3), regardless of models.
We quantify these token-level relations using the same approach as subsubsection 3.2.1. Figure 3(d) shows results across all model layers aggregated from over 24,000 requests. The top 20% of expert pairs account for 40% to 80% of all selections, indicating strong token-level relations. Llama4 exhibits the strongest relation while Deepseek shows the weakest, suggesting future expert selections are most predictable in Llama4.
3.2.3. Prefill-Decode-Level Correlation
@@
To quantify this similarity, we compute Spearman’s ratio ($\rho$) across all model layers, comparing prefill and decode heatmaps. Spearman’s Ratio $\rho$ measures monotonic relationships between variables, ranging from 1-1 (perfect negative correlation) to $1$ (perfect positive correlation). Generally, $|\rho|>0.7$ indicates strong correlation, $0.4<|\rho|\leq 0.7$ indicates moderate correlation, and $|\rho|\leq 0.4$ suggests weak correlation (Schober et al., 2018). The results in Figure 4(e)(f) show that most layers demonstrate strong correlation, while a few show moderate correlation. This makes it possible to predict decode-stage expert selection with prefill-stage data.
3.2.4. System Insights from Temporal Relation
The observed temporal relations in expert selection motivate us to design fine-grained, dynamic strategies on every single unit to reduce its communication with other units. For example, whenever expert weights are read from remote memory, such as remote die DRAM in a multi-chiplet serving system or CXL extension memory in a memory disaggregation system, caching, migration, and prefetching strategies can be deployed to enhance performance.
@@ 3.2.4. System Insights from Temporal Relation
Empirical analysis shows that experts selected more frequently during the prefill stage are also likely to be chosen during the decoding stage. Thus, the expert selection information collected during the prefill phase can serve as an important reference for expert selection prediction at the beginning of the decoding stage, when only a few tokens have been generated and reference data is limited. Furthermore, prefill information can also be leveraged to dynamically adjust expert distribution during the decode stage. This is especially useful in modern serving systems that employ parameter disaggregation, where the prefill and decoding stages may reside on different machines.
@@
Figure 5. Single-expert spatial relation analysis of Llama4 layer 7 shows: (a) non-uniform expert activation distribution; (b) expert selection strongly correlates with task type; (c) expert activation patterns shift significantly when language changes while content remains identical.
3.3. Spatial Relation
@@
3.3.1. Single Expert Activation Imbalance
We examine expert selection frequency at each layer, presenting results for layer 7 of Llama4 in Figure 5. After normalizing selection counts by the average, we observe pronounced skewness where a subset of experts are activated over 16 times more frequently than average. This workload imbalance suggests system designs should duplicate or decentralize frequently used experts.
To investigate selection patterns across different tasks, we analyze all 57 MMLU subjects spanning diverse fields, including biology, history, and math, etc (Hendrycks et al., 2020). Figure 5(b) shows the top 10 most popular experts for each subject. Horizontal bright lines indicate certain experts are consistently activated regardless of subject, while remaining popular experts vary significantly between subjects, demonstrating both overlap and distinction in task-based expert selection.
We further examine task impact using Chinese version of MMLU in MMLU Pro (Wang et al., 2024) with identical questions but different languages. Figure 5(c) reveals distinctly different patterns: although 5-6 experts remain popular across subjects, only two overlap with English MMLU’s most frequently selected experts. This confirms that task characteristics, including language, significantly influence expert selection, enabling task-aware serving systems that optimize expert distribution to balance workloads and reduce data movement.
3.3.2. Expert Pair Co-activation Affinify

### v2

@@
Abstract
Large-scale Mixture of Experts (MoE) Large Language Models (LLMs) have recently become the frontier open weight models, achieving remarkable model capability similar to proprietary ones. But their random expert selection mechanism introduces significant data movement overhead that becomes the dominant bottleneck in multi-unit LLM serving systems.
To understand the patterns underlying this data movement, we conduct comprehensive data-movement-centric profiling across four state-of-the-art large-scale MoE models released in 2025 (200B-1000B) using over 24,000 requests spanning diverse workloads. We perform systematic analysis from both temporal and spatial perspectives and distill six key insights to guide the design of diverse future serving systems. With our insights, we then demonstrate how to improve wafer-scale GPUs as a case study, and show that minor architectural modifications leveraging the insights achieve substantial performance gains, delivering 5.3× and 3.1× average speedups on DeepSeek V3 and Qwen3 respectively. Our work presents the first comprehensive data-centric analysis of large-scale MoE models and a concrete design study using the learned lessons, with profiling traces and simulation framework already open-sourced with >>1k downloads. Our traces and results are publicly available at https://huggingface.co/datasets/core12345/MoE_expert_selection_trace.
@@ I Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across diverse domains, including programming assistance [llm_programming1, llm_programming2], translation [gpt_llm_trans1, lu2024llm_trans2], and chatbots [llm_chatbot1, llm_chatbot2]. Since the beginning of 2025, large-scale Mixture of Experts (MoE) LLMs (200B+ model with 100+ experts) have become the leading models for frontier LLMs [chiang2024chatbot] and the most widely used open weight models.
[…]
Figure 1: MoE LLM models sizes and release dates. Bubble size indicates the number of experts in each layer. Prior studies [zhu2025megascale-infer, tairin2025emoe, skliar2024mixture, chitty2025lexi] provide limited analysis of smaller models from narrow perspectives, while our work presents the first comprehensive analysis of multiple unstudied SOTA models.
@@
If data movement in MoE models were fully unpredictable, it would present significant challenges for deployments on multi-unit systems. From a temporal perspective, the explosive growth in expert combinations would make it impossible to prefetch, cache, or replicate experts in advance. For example, large-scale MoE models like DeepSeek V3 have $C_{256}^{8}=$4,426,165,368$\$ combinations in expert selection. When served with host memory-offload systems, such unpredictability would result in data movement like expert migrations between GPU and host, incurring substantial overhead, as inter-unit communication becomes the primary bottleneck. From a spatial perspective, if expert selection were truly random, it would lead to severe workload imbalance across units. When queries from diverse tasks are served concurrently, the number of queries assigned to each expert would vary dramatically, creating significant workload disparities. Consequently, most units would remain idle and wait for heavily loaded units to finish, resulting in poor hardware resource utilization.
@@
Fortunately, as we later show in the paper, MoE expert selections indeed have predictability that designers can exploit to reduce data movement. To uncover the inherent patterns in MoE models, we conduct a comprehensive data-movement-centric profiling of four state-of-the-art MoE models ranging from 235B to 1000B parameters released in 2025. As highlighted in Figure 1, we profile DeepSeek V3 [liu2024deepseek], Llama4 Maverick [llama4], Qwen3-235B [yang2025qwen3], and Kimi K2 [team2025kimi] across 24,000 requests involving varied tasks, topics, and languages, which consumes >>2000 GPU hours in total. We then collect the expert selection trace of all layers and tokens in each request to create an expert selection database of over 150 GB JSON files. From these extensive traces, we conduct a comprehensive analysis to uncover data movement patterns from both temporal and spatial perspectives, making our findings system-agnostic and applicable to various serving architectures at any scale. We then distill six key insights that serve as a solid foundation to understand MoE data movement and directly inform future MoE LLM serving system design, addressing critical questions that have remained unanswered in the field, such as: Is there a correlation between previously selected experts and those selected later? Are there discernible rules underlying the observed expert selection skewness? Do different tasks tend to activate different experts? Our work represents the first systematic effort to characterize data movement patterns at the scale of up-to 1000B model across a wide range of tasks, providing actionable insights that can guide the design of next-generation MoE serving systems.
@@
  • We propose a comprehensive and systematic data-movement-centric profiling across four latest, large-scale MoE models released in 2025 between 235B and 1000B to uncover the data movement patterns from both temporal and spatial perspectives.
@@
  • We collect over 70,000 expert selection traces across multiple models and datasets, totaling over 150 GB in JSON format, and have open-sourced (>>1k downloads) all traces with our multi-chiplet simulator to facilitate future research on MoE serving systems.
Figure 2: Latency breakdown for different data movement in DeepSeekV3, modeled after various serving configurations, including [ds_sglang_16h20, sglang-deepseek-blogpost, liu2024deepseek].
Figure 3: Inference process of MoE LLMs and the categorization method for our proposed data-centric profiling approach.
II Background
II-A LLM and MoE Model Architecture
Most state-of-the-art LLMs adopt a decoder-only transformer architecture that follows a token-by-token auto-regressive workflow [attention]. As shown in Figure 3(a), after users input queries, the serving process is divided into two stages: the prefill stage and the decode stage. During the prefill stage, all input tokens are processed simultaneously to generate the first output token. The decode stage follows immediately, where tokens are generated sequentially. The generated token from each iteration is appended to the input sequence to produce the next token in the following iteration.
The Mixture of Experts (MoE) mechanism is a state-of-the-art approach to improve LLM performance and has become prevalent among current frontier LLMs [moe-survey]. As shown in Figure 3(b), MoE-based LLMs replace the feed-forward network (FFN) layers in traditional LLMs with MoE layers. In each layer, multiple experts are deployed, and each request is routed to a small subset of the most suitable experts based on a gating mechanism. This innovation enables MoE models to scale model parameters without incurring extra inference overhead, since only a fraction of parameters are activated for each request. However, this mechanism also introduces dynamic randomness, since expert selection is unknown until gating is completed, posing new challenges for serving systems.
II-B Prior MoE Serving Systems
The MoE mechanism constitutes the primary source of data movement overhead in modern serving systems. As illustrated in Figure 2, MoE-related data movement (MoE All-to-All and MoE Weights) dominates the overhead across different DeepSeek V3 serving configurations, accounting for 60%-90% of total latency. To address this overhead, existing research on MoE serving optimization has developed numerous system-level solutions targeting different performance and cost objectives. Edge systems like MoE-Lightning [moelightning] and CoServe [suo2025coserve] employ CPU memory offloading techniques to address GPU memory capacity constraints, while server systems such as Comet [zhang2025comet] and MegaScale-Infer [zhu2025megascale-infer] leverage multi-GPU architectures and address GPU-GPU communications in MoE for higher throughput. Novel hardware architectures like Duplex [yun2024duplex] explore processing-in-memory to accelerate data movement in MoE LLMs.
@@ II-B Prior MoE Serving Systems
Figure 4: Layer-level temporal correlation heatmaps for (a) Deepseek and (b) Qwen, together with (c) statistical results across all layers, reveal a strong correlation between expert selection in layer NN and that in layer $N+1$.
Figure 5: Token-level temporal correlation heatmaps for (a) Deepseek, (b) Llama, and (c) Qwen, together with (d) statistical results across all layers, demonstrate a strong correlation between the expert selection for consecutive tokens $t$ and $t+1$.
Figure 6: Expert activation patterns remain consistent across prefill and decode stages for both (a, b) cross-layer heatmap and (c, d) cross-token heatmaps. Spearman’s ratio quantified in (e, f) shows a strong relation ($\geq 0.7$).
III MoE Profiling and System Insights
In this section, we conduct a data-movement-centric profiling of the expert selection behavior in four state-of-the-art MoE models: Deepseek V3 (671B), Llama4-Maverick-128E (402B), Qwen3-235B (235B), and Kimi K2 (1000B). All results are averaged over more than 24,000 requests.
III-A Categorization Methodology
As in Figure 3, we categorize MoE expert selection profiling results into two categories: temporal and spatial relations.
@@ III-B1 Layer-Level
The heatmaps reveal clear cross-layer correlations with white dots highlighting specific expert pairs with significantly higher selection probabilities across adjacent layers. However, correlation patterns vary across layers within the same model and differ between models due to architectural variations. For instance, patterns between layers 3-4 differ from those between layers 30-31. Qwen3’s notably brighter heatmap indicates stronger cross-layer correlations than Deepseek. Beyond the white dots, there are also consistent bright vertical lines, suggesting certain experts are frequently chosen regardless of previous layer selections. These patterns indicate generally popular experts, analyzed further in Sec. III-C1.
To quantify these relationships, we analyze expert pair frequencies across multiple requests and all model layers. The cumulative percentage of the most frequently selected expert pairs is shown in Figure 4(c). The results demonstrate that the top 20% of expert pairs account for 45%, 68%, 80%, and 55% of all activations in Deepseek, Qwen, Llama4, and Kimi K2, respectively. This reveals strong, model-dependent cross-layer correlations, with Llama4  [^1] Llama4 places dense FFN layers between MoE layers, so we analyze correlations between adjacent MoE layers ($N$ and $N+2$). exhibiting the strongest effect and Deepseek the weakest.
III-B2 Token-Level
@@
Similar to layer-level patterns, cross-token heatmaps exhibit white dots, bright vertical lines, and variation across layers and models, indicating correlations between adjacent tokens. However, token-level relations reveal a common pattern appearing across all models: the bright diagonal line that indicates the tendency to select the same expert across adjacent tokens. This diagonal pattern emerges predominantly in higher layers (17 and 43) but not in lower layers (1 and 3), regardless of models.
We quantify these token-level relations using the same approach as layer-level relations in subsubsection III-B1. As shown in Figure 5(d), the top 20% of expert pairs account for 40% to 80% of all selections, indicating strong token-level relations. Llama4 exhibits the strongest relation while Deepseek shows the weakest, suggesting future expert selections are most predictable in Llama4.
III-B3 Prefill-Decode-Level Correlation
@@
To quantify this similarity, we compute Spearman’s ratio ($\rho$) across all model layers, comparing prefill and decode heatmaps. Spearman’s Ratio $\rho$ measures monotonic relationships between variables, ranging from 1-1 (perfect negative correlation) to $1$ (perfect positive correlation). Generally, $|\rho|>0.7$ indicates strong correlation, $0.4<|\rho|\leq 0.7$ indicates moderate correlation, and $|\rho|\leq 0.4$ suggests weak correlation [spearman]. The results in Figure 6(e)(f) show that most layers demonstrate strong correlation, while a few show moderate correlation. This makes it possible to predict decode-stage expert selection with prefill-stage data.
III-B4 System Insights from Temporal Relation
The observed temporal relations in expert selection motivate us to design fine-grained, dynamic strategies on every single unit to reduce the communications. For example, when expert weights are read from remote memory, such as remote DRAM in multi-chiplet systems, or CXL extension memory in memory-disaggregated systems, caching, migration, and prefetching strategies can be deployed to reduce data movement.
@@ III-B4 System Insights from Temporal Relation
Empirical analysis shows that experts selected more frequently during the prefill stage are also likely to be chosen during the decoding stage. Thus, the expert selection information collected during the prefill phase can serve as an important reference for expert selection prediction at the beginning of the decoding stage, when only a few tokens have been generated and reference data is limited. Furthermore, prefill information can also be used to dynamically adjust expert distribution during the decode stage. This is especially useful in modern serving systems that employ parameter disaggregation, where the prefill and decode stages may reside on different machines.
@@
Figure 7: Single-expert spatial relation analysis of Llama4 layer 7 shows: (a) non-uniform expert activation distribution; (b) expert selection strongly correlates with task type; (c) expert activation patterns shift significantly when language changes while content remains identical.
III-C Spatial Relation
@@
III-C1 Single Expert Activation Imbalance
We examine expert selection frequency at each layer, presenting results for layer 7 of Llama4 in Figure 7. After normalizing selection counts by the average, we observe pronounced skewness where a subset of experts is activated over 16 times more frequently than average. This workload imbalance suggests system designs should duplicate or decentralize frequently used experts.
To investigate selection patterns across different tasks, we analyze all 57 MMLU subjects spanning diverse fields, including biology, history, and math, etc [mmlu]. Figure 7(b) shows the top 10 most popular experts for each subject. Horizontal bright lines indicate certain experts are consistently activated regardless of subject, while remaining popular experts vary significantly between subjects, demonstrating both overlap and distinction in task-based expert selection.
We further examine task impact using Chinese version of MMLU in MMLU Pro [mmlu_pro] with identical questions but different languages. Figure 7(c) reveals distinctly different patterns: although 5-6 experts remain popular across subjects, only two overlap with English MMLU’s most frequently selected experts. This confirms that task characteristics, including language, significantly influence expert selection, enabling task-aware serving systems that optimize expert distribution to balance workloads and reduce data movement.
III-C2 Expert Pair Co-activation Affinify